# Visual Attention Based Cognitive Human–Robot Collaboration for Pedicle Screw Placement in Robot-Assisted Orthopedic Surgery

### v1

@@ III-B1 Eye-Tracking Based Human Attention Recognition
III-B2 Attention-Driven Human-Robot Collaboration
@@ III-C2 Shared Interaction Control on Haptic Device
where $\bm{D}$ and $\bm{K}$ represent the virtual mass, damping, and stiffness matrices to be simulated by the device, $\bm{x}_{d}$ is the equilibrium position (i.e., the desired drilling position), and $\bm{f}_{\text{ext}}$ is the external force applied to the device by the surgeon.
@@
$\bm{u}_{\text{haptic}}=\bm{J}^{\mathsf{T}}(\bm{\theta})(\bm{K}(w)(\bm{x}_{d}-\bm{x})-\bm{D}\dot{\bm{x}}+\bm{f}_{\text{fdbk}})+\bm{g}(\bm{\theta})$ (9)
@@ IV Experiments
In order to evaluate the effectiveness of the proposed framework, experiments were conducted using an orthopedic surgical robot system specifically designed for pedicle screw insertion tasks (see Fig. 5). The robot featured a UR5 cobot from Universal Robots, equipped with a custom bone drill end-effector. The end-effector is outfitted with an ATI mini40 f/t sensor to monitor drilling forces and an Intel RealSense D405 RGB-D camera to oversee the surgical site. Users interacted with the robot using a Force Dimension Omega.3 haptic device. The AR scene was developed on the Unity engine with MRTK3 and was deployed on Microsoft HoloLens 2 AR-HMD. The software system was developed and integrated using the ROS 2 [34], MoveIt 2 [35] and ros2_control [36].
@@ V Conclusions
In this paper, we proposed a cognitive human-robot collaboration framework for robot-assisted orthopedic surgery, based on AR-haptic interface, surgeon visual attention model and shared control. This work improves the HRI experience of orthopedic surgical robots and has the potential to increase surgical efficiency and safety. Moreover, the proposed framework provides insights into the design of collaborative robotic systems in safety-critical tasks. Future works will be devoted to the validation of the developed system in clinical trials.

### v2

@@ III-B1 Eye-Tracking Based Human Attention Recognition
III-B2 Attention-Driven Human–Robot Collaboration
@@ III-C2 Shared Interaction Control on Haptic Device
where $\bm{D}$ and $\bm{K}$ represent the virtual damping and stiffness matrices to be simulated by the device, $\bm{x}_{d}$ is the equilibrium position (i.e., the desired drilling position), and $\bm{f}_{\text{ext}}$ is the external force applied to the device by the surgeon.
@@
$\bm{u}_{\text{haptic}}=\bm{J}^{\mathsf{T}}(\bm{\theta})(\bm{K}(w)(\bm{x}_{d}-\bm{x})-\bm{D}\dot{\bm{x}}+\bm{f}_{\text{fdbk}})+\bm{g}(\bm{\theta}),$ (9)
@@ IV Experiments
[…]
@@ V Conclusions
In this paper, we proposed a cognitive human–robot collaboration framework for robot-assisted orthopedic surgery, based on AR-haptic interface, surgeon visual attention model and shared control. This work improves the HRI experience of orthopedic surgical robots and has the potential to increase surgical efficiency and safety. Moreover, the proposed framework provides insights into the design of collaborative robotic systems in safety-critical tasks. Future works will be devoted to the validation of the developed system in clinical trials.
Acknowledgment
The authors want to thank Yu Chen for the initial prototype, and thank all subjects for taking part in the human study.